OpenStudy (anonymous):

Which angle (in radians) corresponds to 2/3 of a circle?

OpenStudy (campbell_st):

a circle is 2 pi radians so find 2/3 of 2 pi

OpenStudy (anonymous):

so would it be 7pi/8 or 240 degrees

OpenStudy (campbell_st):

\[\frac{2}{3} \times 2 \pi = \frac{4\pi}{3}\]its actually